OpenStudy (anonymous):

Write your answer as a power 2. Show your work and explain your steps. Computer capacity is often measured in bits and bytes. A bit is the smallest unit, which is a 1 or 0, in the computer’s memory. A byte is 2^3 bits. A megabyte (MB) is 2^20 bytes. How man bits are in a megabyte?

OpenStudy (anonymous):

well 1 byte = 2^3 2^20 = x cross multiply so 2^20*2^3 = x 4^23 = x you want it in bits, so you might want to turn 4 in to 2 i think so factor it out which will be 2(2^23) -- i think this is right correct it me if i am wrong